a weekend trip package to the beach includes a two night hotel stay and four meals. a week trip includes a seven night hotel stay and ten meals. the 2 night stay has a price of $439.94 and the seven night stay has a price of $1419.83. find the cost of one hotel stay and one meal
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

2h+4m = 439.94 multiply all times 7:
14h+28m = 3079.58 (1)
--------------------------------------------------------
7h+10m = 1419.83
multiply all times -2:
-14h+-20m = -2839.65 (2)
----------------------------------------------------------
Add (1) and (2)
14h+28m = 3079.58
+
-14h+-20m = -2839.65
-------------------
0h + 8m = 239.93
8m = 239.93
m = 239.93/8 = 29.99125, i'll round it off to 30.
-----------------------------------------------------------
now that we know m we can solve for h:
2h+4m = 439.94
2h+4(30) = 439.94
2h+120 = 439.94
2h = 319.94
h = 159.97
So there you have it, the hotel was 160 (159.97 rounded off) and the meals were 30.

Answer by MathTherapy(10552)


a weekend trip package to the beach includes a two night hotel stay and four meals. a week trip includes a seven night hotel stay and ten meals. the 2 night stay has a price of $439.94 and the seven night stay has a price of $1419.83. find the cost of one hotel stay and one meal
Let one night be H, and one meal, M
Then we get: 2H + 4M = 439.94____2(H + 2M) = 2(219.97)____H + 2M = 219.97____H = 219.97 - 2M --------- eq (i)
7H + 10M = 1,419.83 ------- eq (ii)
7(219.97 - 2M) + 10M = 1,419.83 -------- Substituting 219.97 - 2M for H in eq (ii)
1,539.79 - 14M + 10M = 1,419.83
- 4M = - 119.96
M, or cost of one meal =
Cost to stay 1 night =